# Resonance Reduction Against Adversarial Attacks

### v1

@@ VI-C2 Parameter Analysis
Parameters that specify an eigenspectrum optimization problem on an auxiliary graph is similar to those of network graph optimization. Since the auxiliary graph edges and inter-graph edges are assumed to have non-negative weights, we do not consider the minimum weight constraint ($w^{min}$) parameter in this case. However, in addition to the network graph optimization parameters, we must consider the effects of the following parameters associated with auxiliary graphs: the auxiliary connectivity type (mirrored or complete), the weights resource multiplier $r_{m}$, and the auxiliary damping factor ${\widetilde{\gamma}}$. We defer the analysis of the auxiliary damping factor to Section VI-C4 and use a constant auxiliary damping factor of ${\widetilde{\gamma}}=10^{-6}$ throughout the parameter analysis.
We analyze the effects of these parameters on the percentage reduction of objective value that can be achieved via the auxiliary graph optimization, hence the relative decrease in the vulnerability of the graph using the AGO method. We start with the same set of parameter values with the addition of $r_{m}=5$, and generate problem instances where we vary one parameter and keep the rest constant. We solve for each problem instance and compute the percentage reductions in objective as %$\%d_{{\widetilde{J}}}=\frac{|{J}^{0}-{\widetilde{J}}^{*}|}{{J}^{0}}\times 100$, which are plotted against the varying parameter values in Figure 16. Note that for the problem instances where the main network graph is a RIG, we provide two sets of results achieved with a mirrored auxiliary graph and a complete auxiliary graph.
@@
For all parameters, effects are similar to those on the network graph optimization. However, even for the parameter values for which the NGO was less effective, the AGO method can achieve larger decreases in the objective, which makes the approach less sensitive to the choice of the parameters. The same insensitivity is observed to the weight resource multiplier parameter. For the instances where the network graph was incomplete, some of the optimizations of the mirrored auxiliary graph failed to converge in the maximum number of iterations considered, which is indicated by a $0\%$ decrease in the plots.
VI-C3 Demonstration of the Effectiveness of AGO
To demonstrate the overall effectiveness of eigenspectrum optimization on the auxiliary graph in reducing the network vulnerability, we solve the optimization problem for RCGs and RIGs and show that significant decrease in objective values can be achieved. We use the same problem instances generated for the NGO, with $r_{m}=5$ and ${\widetilde{\gamma}}=10^{-6}$ and using complete auxiliary graphs. To demonstrate the effects of attaching an arbitrary auxiliary graph and the optimization of this auxiliary graph separately, we provide the average and the standard deviation of the percentage decrease in the objective calculated as (1) %$\%d_{{\widetilde{J}}}=\frac{|{J}^{0}-{\widetilde{J}}^{0}|}{{J}^{0}}\times 100$ (decrease achieved by going from network configuration ${G}$ to ${G}+{\widetilde{G}}$), and (2) %$\%d_{{\widetilde{J}}}=\frac{|{J}^{0}-{\widetilde{J}}^{*}|}{{J}^{0}}\times 100$ (decrease achieved by going from network configuration ${G}$ to ${G}+{\widetilde{G}}^{*}$) across the problem instances featuring complete and incomplete main network graphs are provided in Figure 17.
@@ VI-C4 Effect of the Auxiliary Damping and Auxiliary Damping Optimization
VII Conclusion
@@
In this paper, we developed the notion of vulnerability of a network with second order signal dynamics under adversarial forcing that obeys a known stochastic model. To minimize the network vulnerability, we proposed two methods that optimize the network structure: i. The Network Graph Optimization (NGO) method provides an optimal set of network edge weights under the condition that the edge weights can be directly manipulated, and, ii. The Auxiliary Graph Optimization (AGO) method allows us to design an auxiliary network that can be attached to the main network with the purpose of minimizing the vulnerability, when the main network edge weights cannot be adjusted directly. We conducted numerical experiments to analyze the two methods in detail.
@@
Currently, the notion of vulnerability and the optimization problems posed in this work depend on a linear model of the signal dynamics and a specific stochastic model of adversarial forcing. While the adaptation of some aspects of the model to other setting (e.g., a different stochastic model of adversarial forcing) can be straight-forward re-derivation of the objective functions, a more general formulation that encompasses more complicated signal models, forcing models, and potentially nonlinear signal dynamics, is within the scope of future work. The optimization formulations presented in this paper lead to generally non-convex problems which are in turn solved by gradient based solvers. A formal analysis of the optimization landscape would be necessary to provide any guarantees on the quality of the solution being returned and is within the scope of future work.
@@
The current optimization problem is formulated as a centralized one that assumes complete knowledge of the network graph edge weights. A potential future work involves the development of a distributed optimization scheme in which each vertex would use information about its local subgraph and would only adjusts weights on its incident edges in order to optimize the network. A distributed method would allow the approach to scale to larger networks and generalize to settings where global information regarding the network may not be available due to privacy restrictions. Another front of exploration is the application of the eigenspectrum optimization principle to real-world dynamics over physical networks, such as electrical grids, robot networks and social networks, specifically focusing on the translation of the constraints on a real network to optimization constraints and the demonstration of resonance mitigation in such realistic scenarios.

### v2

@@ VI-C2 Parameter Analysis
Parameters that specify an spectrum optimization problem on an auxiliary graph is similar to those of network graph optimization. Since the auxiliary graph edges and inter-graph edges are assumed to have non-negative weights, we do not consider the minimum weight constraint ($w^{min}$) parameter in this case. However, in addition to the network graph optimization parameters, we must consider the effects of the following parameters associated with auxiliary graphs: the auxiliary connectivity type (mirrored or complete), the weights resource multiplier $r_{m}$, and the auxiliary damping factor ${\widetilde{\gamma}}$. We defer the analysis of the auxiliary damping factor to Section VI-C4 and use a constant auxiliary damping factor of ${\widetilde{\gamma}}=10^{-6}$ throughout the parameter analysis.
We analyze the effects of these parameters on the percentage reduction of objective value that can be achieved via the auxiliary graph optimization, hence the relative decrease in the vulnerability of the graph using the Auxiliary Graph Optimization method. We start with the same set of parameter values with the addition of $r_{m}=5$, and generate problem instances where we vary one parameter and keep the rest constant. We solve for each problem instance and compute the percentage reductions in objective as %$\%d_{{\widetilde{J}}}=\frac{|{J}^{0}-{\widetilde{J}}^{*}|}{{J}^{0}}\times 100$, which are plotted against the varying parameter values in Figure 13. Note that for the problem instances where the main network graph is a RIG, we provide two sets of results achieved with a mirrored auxiliary graph and a complete auxiliary graph.
@@
For all parameters, effects are similar to those on the network graph optimization. However, even for the parameter values for which the network graph optimization was less effective, the Auxiliary Graph Optimization method can achieve larger decreases in the objective, which makes the approach less sensitive to the choice of the parameters. The same insensitivity is observed to the weight resource multiplier parameter. For the instances where the network graph was incomplete, some of the optimizations of the mirrored auxiliary graph failed to converge in the maximum number of iterations considered, which is indicated by a $0\%$ decrease in the plots.
VI-C3 Demonstration of the Effectiveness of Auxiliary Graph Optimization
To demonstrate the overall effectiveness of spectrum optimization on the auxiliary graph in reducing the network vulnerability, we solve the optimization problem for RCGs and RIGs and show that significant decrease in objective values can be achieved. We use the same problem instances generated for the network graph optimization, with $r_{m}=5$ and ${\widetilde{\gamma}}=10^{-6}$ and using complete auxiliary graphs. To demonstrate the effects of attaching an arbitrary auxiliary graph and the optimization of this auxiliary graph separately, we provide the average and the standard deviation of the percentage decrease in the objective calculated as (1) %$\%d_{{\widetilde{J}}}=\frac{|{J}^{0}-{\widetilde{J}}^{0}|}{{J}^{0}}\times 100$ (decrease achieved by going from network configuration ${G}$ to ${G}+{\widetilde{G}}$), and (2) %$\%d_{{\widetilde{J}}}=\frac{|{J}^{0}-{\widetilde{J}}^{*}|}{{J}^{0}}\times 100$ (decrease achieved by going from network configuration ${G}$ to ${G}+{\widetilde{G}}^{*}$) across the problem instances featuring complete and incomplete main network graphs are provided in Figure 14.
@@ VI-C4 Effect of the Auxiliary Damping and Auxiliary Damping Optimization
Figure 16: Example of eigenvalue spectrum of a complete graph before and after optimization represented as histograms of the eigenvalues of the networks stiffness matrix, $L+\varepsilon I$. The objective value significantly decreased from 1.378 to 0.3778, yielding a 72.58% decrease as a result of the network graph optimization on a 100-node RCG. Qualitatively, as a result of the network graph optimization, the eigenvalue spectrum has become flatter.
Figure 17: Example of eigenvalue spectrum of a representative Facebook Social subgraph before and after optimization represented as histograms of the eigenvalues of the networks stiffness matrix, $L+\varepsilon I$. The objective value significantly decreased from 6.868 to 2.467 as a result of the network graph optimization on the 173-node Facebook social subgraph, corresponding to a 64.089% decrease in the objective value. As a result of the optimization, the eigenvalue spectrum has become smoother and flatter.
Figure 18: An embedding of an example Facebook Social Subgraph with 173 vertices. The colors on the edges indicate weights after optimization.
VII Conclusion and Discussions
@@
In this paper, we developed the notion of vulnerability of a network with second order signal dynamics under adversarial forcing that obeys a known stochastic model. To minimize the network vulnerability, we proposed two methods that optimize the network structure: i. The Network Graph Optimization method provides an optimal set of network edge weights under the condition that the edge weights can be directly manipulated, and, ii. The Auxiliary Graph Optimization method allows us to design an auxiliary network that can be attached to the main network with the purpose of minimizing the vulnerability, when the main network edge weights cannot be adjusted directly. We conducted numerical experiments to analyze the two methods in detail.
@@
Currently, the notion of vulnerability and the optimization problems posed in this work depend on a linear model of the signal dynamics and a specific stochastic model of adversarial forcing. While the adaptation of some aspects of the model to other setting (e.g., a different stochastic model of adversarial forcing) can be straight-forward re-derivation of the objective functions, a more general formulation that encompasses more complicated signal models, forcing models, and potentially nonlinear signal dynamics, is within the scope of future work. The optimization formulations presented in this paper lead to generally non-convex problems which are in turn solved by gradient based solvers. While we do show convexity (Proposition LABEL:prop:convexity) of the objective function of the Network Graph Optimization problem under the assumption that the parameter $h$ is large, a more general analysis of the optimization landscape for finite values of $h$ would be necessary to provide guarantees on the quality of the solution being returned, both for the Network Graph Optimization as well as the Auxiliary Graph Optimization problems. Such analyses are within the scope of future work.
@@
The current optimization problem is formulated as a centralized one that assumes complete knowledge of the network graph edge weights. A potential future work involves the development of a distributed optimization scheme in which each vertex would use information about its local subgraph and would only adjusts weights on its incident edges in order to optimize the network. A distributed method would allow the approach to scale to larger networks and generalize to settings where global information regarding the network may not be available due to privacy restrictions. In future we will work towards implementing the proposed methods on real-world, physical networks such as electrical grids, robot networks and social networks.